The boat has an anchor on a 90ft line. you take the boat out to the middle of the lake on a windy day. you drop the anchor and let the wind push the boat until the anchor line it tight. your GPS tells you that the boat has moved 82 feet. assuming the bottom of the lake is flat, what is the depth of the lake? simplify your answer, then round it to the nearest foot.
Answer by stanbon(75887)

The boat has an anchor on a 90ft line. you take the boat out to the middle of the lake on a windy day. you drop the anchor and let the wind push the boat until the anchor line it tight. your GPS tells you that the boat has moved 82 feet. assuming the bottom of the lake is flat, what is the depth of the lake? simplify your answer, then round it to the nearest foot.
------
Sketch the right triangle including rope = 90 ft and base = 82
-----
rope is the hypotenuse
depth is the 3rd side of the triangle
----
d^2 + 82^2 = 90^2
d^2 = 8100-6724 = 1376
depth = 37.09 ft
------
Cheers,
Stan H.
----------